Skip to content

Bugfix and code cleanup for Ticket #1095.

liufeiyang requested to merge yang/VVCSoftware_VTM:Bugfix_Ticket1095 into master

Rate control in current VTM does not work. Issues are pointed out in Ticket #1095. This merge request is provided to fix them.

  1. Fix the wrong clipping of alpha and beta for intra frames.
  2. Fix the variable of number of CTUs left in rate allocation for intra frames.
  3. Fix the updating of alpha and beta for intra frames.
  4. Fix the missing chroma distortion weight for CTU level rate control.
  5. Code cleanup.

Merge request reports